# Enhancing ASR Performance through OCR Word Frequency Analysis: Theoretical Foundations

### v4

@@ Abstract
As the interest in large language models grows, the importance of accuracy in automatic speech recognition has become more pronounced. This is particularly true for lectures that include specialized terminology, where the success rate of traditional ASR models tends to be low, posing a challenging problem. A method to improve ASR performance for specialized terminology using the word frequency difference approach has been proposed. Through experiments and data analysis, we investigated whether this proposal effectively addresses this issue. In addition, we introduced the power law as the theoretical foundation for the relative frequency methodology mentioned in this approach.
Figure 1: Figure 1. (a) Graph of RF using the existing method [ref3], with log(RF rank) on the x-axis and log(RF) on the y-axis. (b) Graph of RF for data processed using Method 1 in addition to the existing method. (c) Graph of RF for data processed using both Method 1 and Method 2 in addition to the existing method.
@@ I Introduction
Recent advancements in research and development of conversational AI and large language models (LLMs) have been notably vigorous. These technological advancements extend their potential applications not only to everyday conversations but also to specialized fields such as medicine, law, engineering, and education. To facilitate this, Automatic Speech Recognition (ASR) systems must be capable of accurately recognizing and understanding specialized terminology within these fields. Consequently, there is ongoing research to enhance the recognition rates of specialized terminology. Ma’s study implemented a method in which lip movement analysis provides clues to enable ASR systems to recognize unfamiliar words [ref1]. This represents an innovative attempt to utilize visual information to compensate for unclear pronunciation in audio signals. Furthermore, Guo’s research involved an approach to correct inaccuracies in ASR outputs by incorporating spelling-correction efforts [ref2]. Guo proposed a method that assigns weights according to spelling rules to the n-best sequences obtained from the ASR model and selects the sequence with the highest score.
@@
In this study, we experimented with Jung’s method. We meticulously dissected and analyzed the method proposed by Jung. The results identified two problems that could be effectively resolved through the development of this method. Additionally, the theoretical background of the Relative Frequency (RF) methodology was validated through the power law.
@@ III-A2 Data Analysis of the Experiment on the Existing Method
This confirmed that the existing method decreases the reliability and accuracy of RF values by uniformly assigning high RF values to words not found in the LTD.
@@ III-A3 Experiment on the Improved Method 1
When applying the improved data processing method proposed in [ref4], enhanced RF values were observed. The datasets from both LTD and OCRd were merged based on the intersection of words in the ’word’ column of the latter. This process excludes words from the LTD that are not extracted by the OCR. When a word extracted by OCR is not in LTD, its count value is replaced by the minimum count value from the merged dataset. The resulting graph is shown in Fig.1b.
@@ IV Conclusion
In this paper, we demonstrated the validity of the method proposed in [ref4] as an improvement over the method in [ref3]. Specifically, we showed through experimentation that the approach of [ref3] has drawbacks in three aspects and that the method in [4] can resolve these shortcomings. Notably, the improved method has a convincing theoretical foundation based on the power law.

### v6

@@ Abstract
As the interest in large language models grows, the importance of accuracy in automatic speech recognition has become more pronounced. This is especially true for lectures that include specialized terminology. In such cases, the success rate of traditional ASR models tends to be low, presenting a significant challenge. A method using the word frequency difference approach has been proposed to improve ASR performance for specialized terminology. We investigated this proposal through experiments and data analysis to determine if it effectively addresses the issue. In addition, we introduced the power law as the theoretical foundation for the relative frequency methodology mentioned in this approach.
Index Terms:
Automatic Speech Recognition, Optical Character Recognition, Power Law, Analysis
@@ I Introduction
Recent advancements in conversational AI and large language model (LLM) research have been notably vigorous. These advancements extend their applications to everyday conversations and specialized fields like medicine, law, engineering, and education. To facilitate this, Automatic Speech Recognition (ASR) systems must be capable of accurately recognizing and understanding specialized terminology within these fields. Consequently, there is ongoing research to enhance the recognition rates of specialized terminology. Ma’s study implemented a method in which lip movement analysis provides clues to enable ASR systems to recognize unfamiliar words [ref1]. This represents an innovative attempt to utilize visual information to compensate for unclear pronunciation in audio signals. Furthermore, Guo’s research involved an approach to correct inaccuracies in ASR outputs by incorporating spelling-correction efforts [ref2]. Guo proposed a method that assigns weights according to spelling rules to the n-best sequences obtained from the ASR model and selects the sequence with the highest score.
@@
In this study, we experimented with Jung’s method, dissecting and analyzing it in detail. The results identified two problems that could be effectively resolved through the development of this method. Additionally, the theoretical background of the Relative Frequency (RF) methodology was validated through the power law.
Figure 1: (a) Graph of RF using the existing method [ref3], with log(RF rank) on the x-axis and log(RF) on the y-axis. (b) Graph of RF for data processed using Method 1 in addition to the existing method. (c) Graph of RF for data processed using both Method 1 and Method 2 in addition to the existing method.
@@ III Analysis and Experiments
This section introduces the methodologies and experiments designed to address discrepancies between text extracted by OCR and a large-scale language corpus. We analyze the issues arising when OCR-extracted words are not found in the large text database (LTD), comparing the performance of existing and improved methods.
@@ III-A2 Data Analysis of the Experiment on the Existing Method
This confirmed that the existing method reduces RF reliability and accuracy by uniformly assigning high RF values to words not found in the LTD.
@@ III-A3 Experiment on the Improved Method 1
When applying the improved data processing method proposed in [ref4], enhanced RF values were observed. The datasets from both LTD and OCRd were merged based on the intersection of words in the ’word’ column of the latter. This process excludes words from the LTD that are not extracted by the OCR. When a word extracted by OCR is not in LTD, its count value is replaced by the minimum count value from the merged dataset. The resulting graph appears in Fig.1b.
@@ IV Conclusion
In this paper, we demonstrated the validity of the proposed method as an improvement over the previous approach. Specifically, we showed through experimentation that the approach of has drawbacks in three aspects and that the method in can resolve these shortcomings. Notably, the improved method has a convincing theoretical foundation based on the power law.